the length of a rectangular parking lot is 10 meters less than twice its width,and the perimeter is 412 meters, find the length of the parking lot.
Answer by chriswen(106)

Let x m be the width of the parking lot.
Let 2x-10 m be the length of the parking lot.
...
P=2(l+w)
412=2(2x-10+x)
412=2(3x-10)
412=6x-20
6x=412+20
6x=432
x=432/6
x=72
2x-10=134
...
Therefore, the length of the parking lot is 134 m.